WRITE THE CODE IN VERILOG: Instead of using Registers, USE D FLIP FLOPS and a clock. Include the logic for a resetA sequential circuit with three D flip-flops A, B, and C, a trigger x, and an output z1, and zo. On this state machine diagraD Flip Flop Example module DEF (clk, in, out); parameter n = 1; input cik; input (n-1:0) in; output (n-1:0] out; reg (n-1:01
Answer #1

Stale díograin olol Joo Toot nay Sale apet react * Be Given 3 D-Bep flops named A, QC input Outpute ZI, 20. let D-flip flops21-EM (3,4,5,8,9) D_A = Em (3.4) D-B = Em (3,9) Đọc = f ( $8) 22= Em ( 2,3, 819) * k-map Simplification 5 O OLH10 AB too ol XCircuit diagram using D-flipflops DA A FE2 - clocks Verilog module for state Machine Averillog module for D-flip flop Code moendmodule // verlog module-top level module ber State machine I using three D-feipflops modele State - machine ll Input port